OpenStudy (litzyceron):

Stephanie made 36 muffins. She made 24 apple muffins and the rest were banana. What is the ratio of apple muffins to banana muffins? A. 2 to 3 B. 12 to 36 C. 24 to 36 D. 2 to 1

OpenStudy (mathmale):

1. determine how many banana muffins were made. 2. Now try to answer the posted question.

OpenStudy (calculusxy):

36 muffins in total 24 apple muffins x banana muffins \(x + 24 = 36\) solve for x to get the # of banana muffins

OpenStudy (litzyceron):

but how do i find the ratio?

OpenStudy (mathmale):

24 apple ?? banana find this ratio: number of apple muffins ------------------------ number of banana muffins

OpenStudy (calculusxy):

you need to find the number of banana muffins first and then put it as a ratio with the # of apple muffins (which is already given to us as 24 apple muffins)

OpenStudy (litzyceron):

so it would be 34 to 12 and that is not an option

OpenStudy (calculusxy):

did you solve for x in \(x + 24 = 36\)

OpenStudy (mathmale):

Please, litzyceron, read through this entire discussion and do the work suggested. If you understand that material, y ou'll likely get the correct answer in the end. No short cuts.

OpenStudy (mathmale):

We've got to know how many banana muffins were made. calculusxy and I have both discussed with you the question of how to find that number.

OpenStudy (mathmale):

This problem, like many in math, involves several steps. You seem to be so anxious to finish that you're skipping some of the steps. Please go back to the beginning of this conversation and read it through. Please find the number of bananas.

OpenStudy (mathmale):

calculusxy wrote the following to you: did you solve for x in x+24=36 Please solve for x. That's the # of bananas.